### v1

# Tensor invariants of finite and classical groups

Abstract.
Given a faithful representation $W$ of a finite group $G$, we classify when the invariant subring of the tensor algebra $T(W)$ is finitely generated/noetherian as a twisted commutative algebra (tca). For a representation $W$ of a classical group $G_{\mathbf{Z}}$, we show that the invariant subring $T(W_{k})^{G_{k}}$ is a finitely generated tca for $k$ algebraically closed of sufficiently large characteristics, assuming $W$ admits a good filtration over $\mathbf{Z}$. Finally, we introduce a categorical variant of the Gelfand–Kirillov dimension and compute its value to be $\binom{n+1}{2}$ for $T(\mathbf{C}^{n})$ as a tca.
1. Introduction
An overarching goal of this paper is to demonstrate how Schur–Weyl duality enables us to answer questions in noncommutative ring theory using standard methods from commutative algebra, even in positive characteristic. We specifically address the finite generation problem for tensor invariants by relating it to Weyl’s polarization technique using the Schur functor. Throughout this paper, we let $k$ be an infinite field of characteristic $p\geq 0$.
1.1. Noncommutative invariants.
Given a finite group $G$ acting on a finite-dimensional vector space $W$, it is well known that the invariant subring $\operatorname{Sym}(W)^{G}$ is finitely generated. The analogous question for the tensor algebra $T(W)$ is quite subtle with a rich history. Wolf [Wol36] showed that the ring of noncommutative symmetric polynomials in $\geq 2$ variables is not finitely generated. Almost half a century later, Dicks–Formanek [DF83] proved this almost always holds: the subring of tensor invariants $T(W)^{G}$ is a finitely generated $k$-algebra if and only if $G$ acts via a finite cyclic group consisting of scalar matrices.
Kharčenko recognized the significance of the commuting actions of the symmetric groups $\mathfrak{S}_{n}$ on $T(W)_{n}$ to this question. He [Kha84] asked whether $T(W)^{G}$ admits a finite generating set when incorporating this action, noting that this holds for soluble groups $G$. Koryukin [Kor84] answered Kharčenko’s question affirmatively for linearly reductive groups. To that end, he introduced the notion of an “S-algebra” and proved a noetherian result for $T(W)$ as an S-algebra. From this, he deduced that $T(W)^{G}$ is also a finitely generated S-algebra using the Reynolds operator. Over the past decade, Sam–Snowden [SS12, SS17] and others have independently studied S-algebras under the name “twisted commutative algebras” (tcas), and we adopt this terminology exclusively. See Section 2.2 for definitions.
1.2. Our results.
We settle all remaining cases of the finite generation problem for finite group tensor invariants in the negative while further strengthening Koryukin’s positive result.
Theorem 1.1.
Assume $G\subset GL(W)$ is a finite group. As a twisted commutative algebra, the invariant subring $T(W)^{G}$
  • is noetherian and generated by elements of degree #$\leq\#G$, if $p\nmid\#G$ or $p=0$; and
  • is not generated in finitely many degrees, if $p\mid\#G$.
The infinite generation of the invariant ring for the permutation action of $\mathfrak{S}_{p}$ on $T(k^{p})$ as a tca when $p>0$ was obtained only recently in [BDDK23] (see Example 3.6). Our contribution to the first part is twofold: we establish a module-theoretic variant of Koryukin’s theorem where, in our terminology, he only shows that $T(W)^{G}$ is weakly noetherian; and we make Koryukin’s result effective.
Our key idea is to bridge the study of tensor invariants with the invariants of certain commutative—albeit infinite-dimensional—algebras via Schur–Weyl duality and then invoke results from classical invariant theory. We recall relevant results about the Schur functor in Section 2. In Section 3, we prove Theorem 1.1 by transporting results of Richman [Ric96] about the failure of Weyl’s polarization theorem for finite groups in the modular setting and of Fleischmann extending Noether’s bound to small characteristics [Fle00] to the noncommutative case.
While the first part of our main theorem without the generation bound extends easily to infinite linearly reductive groups, the second part does not immediately generalize to infinite reductive groups. We partially address this with our next result. Let $G$ be a classical group and $W$ be a finite $\mathbf{Z}$-module with a linear action of $G$ affording a universal good filtration; see Section 4.1. The base change $W_{k}=W\otimes k$ is a representation of $G_{k}=G\times_{\mathbf{Z}}k$. By carrying over results of Derksen–Makam [DM21] through our bridge, we obtain:
Theorem 1.2.
There exists $N>0$ such that for all algebraically closed fields $k$ of characteristic $p>N$, the tca $T(W_{k})^{G_{k}}$ is finitely generated.
In Example 4.2, we show that the conclusion of the above theorem can fail for small positive values of $p$. The structure of the invariant rings occurring in Theorem 1.2 is of great interest, as determining whether or not they are noetherian is a major open problem in the field of representation stability (see [Gan24] and references therein). We thus prove some auxiliary results about their structure (Theorem 4.3).
In the final section, we introduce the Gelfand–Kirillov dimension for finitely generated algebra objects in a locally finite length symmetric monoidal category and compute it for some tcas. Notably, in characteristic zero, $\operatorname{GKdim}$ of polynomial tcas generated in degree one agrees with the Krull–Gabriel dimension of their module category as computed by Nagpal–Sam–Snowden [SS19, Section 3.5]. It would be interesting to prove a general result along these lines.
1.3. Related work
Tambour and Teranishi independently pioneered the use of Schur–Weyl duality to study tensor invariants, albeit only in characteristic zero. In [Ter88], Teranishi explains how to explicitly obtain generators of $T(W)^{G}$ as a tca using generators of $\operatorname{Sym}(W^{\oplus\dim W})^{G}$, a method made possible by combining Schur–Weyl duality with Weyl’s polarization theorem. He [Ter91] also provides a novel method to derive generators of $T(W)^{G}$ from that of $\operatorname{Sym}(W\oplus{\textstyle\bigwedge}^{2}W)^{G}$. In his thesis [Tam92], Tambour reproves Koryukin’s theorem, obtains degree bounds for the invariant ring, and introduces and proves rationality results for a certain enhanced Hilbert series by using classical results from invariant theory like Noether’s bound and Molien’s formula.
Twisted (commutative) algebras, even infinitely generated ones, arise naturally when studying the topology of configuration spaces of points/discs on a plane (see for example, [AM24, Waw23] and references therein). We hope the method developed in Section 5 can be used to provide interesting invariants of topological spaces.
2. Setup
Throughout this paper, a partition $\lambda=(\lambda_{1},\lambda_{2},\ldots)$ is an infinite sequence of integers with $\lambda_{i}\geq\lambda_{i+1}$ for all $i\geq 1$ and $\lambda_{j}=0$ for $j\gg 0$. The size of the partition $|\lambda|=\sum_{i}\lambda_{i}$. A partition is $p$-restricted if $\lambda_{i}-\lambda_{i+1}<p$ for all $i\geq 1$. We also let $\mathbf{V}$ be the infinite-dimensional $k$-vector space with basis $e_{1},e_{2},\ldots$ and $\mathbf{GL}$ be the group of $k$-linear automorphisms of $\mathbf{V}$.
2.1. Symmetric monoidal categories
For definitions about symmetric monoidal categories and functors, we refer the reader to [Rei16, Chapter 2]. We first introduce the two symmetric monoidal categories that are central to this paper.
  1. (1)
    $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL}):$ A representation of $\mathbf{GL}$ is polynomial if it occurs as a subquotient of direct sums of $\mathbf{V}^{\otimes n}$ with $n$ allowed to vary. We let $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$ denote the category of polynomial representations of $\mathbf{GL}$; it is a symmetric monoidal category with the usual tensor product of representations (denoted $\otimes$). The simple objects in this category are indexed by arbitrary partitions $\lambda$ and denoted $L_{\lambda}(\mathbf{V})$.
  2. (2)
    $\operatorname{Rep}(\mathfrak{S}_{*})$: An $\mathfrak{S}_{*}$-representation is a sequence $\{V_{n}\}_{n\in\mathbf{N}}$ where each $V_{n}$ is a representation of the symmetric group $\mathfrak{S}_{n}$ over $k$. We let $\operatorname{Rep}(\mathfrak{S}_{*})$ be the category of $\mathfrak{S}_{*}$-representations. Given objects $V\coloneqq\{V_{n}\}$ and $W\coloneqq\{W_{n}\}$, we define
    $(V\boxtimes W)_{n}=\bigoplus_{i=0}^{n}\operatorname{Ind}_{\mathfrak{S}_{i}\times\mathfrak{S}_{n-i}}^{\mathfrak{S}_{n}}V_{i}\boxtimes W_{n-i}.$
    This endows $\operatorname{Rep}(\mathfrak{S}_{*})$ with the structure of a symmetric monoidal category. The simple objects in this category are indexed by $p$-restricted partitions $\lambda$ and denoted $D^{\lambda}$.
The above two categories have a canonical $\mathbf{N}$-grading with $L_{\lambda}(\mathbf{V})$ (and $D^{\lambda}$) homogeneous of degree $|\lambda|$. The monoidal structure preserves this grading on both categories.
2.2. Commutative algebra objects
Definition 2.1.
A $\mathbf{GL}$-algebra is a commutative algebra object in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$. Concretely, a $\mathbf{GL}$-algebra $S$ is a commutative $k$-algebra $S$ with an action of $\mathbf{GL}$ by $k$-algebra automorphisms under which $S$ forms a polynomial representation of $\mathbf{GL}$. ∎
The prototypical example of a $\mathbf{GL}$-algebra is the polynomial ring $\operatorname{Sym}(U)$ where $U$ is a polynomial representation of $\mathbf{GL}$. In this paper, we will mostly be concerned with the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})$ where $W$ is a finite-dimensional $k$-vector space with trivial $\mathbf{GL}$-action.
Definition 2.2.
A twisted commutative algebra (tca) is a commutative algebra object in $\operatorname{Rep}(\mathfrak{S}_{*})$. Concretely, a tca is an associative, unital, $\mathbf{N}$-graded $k$-algebra $A=\bigoplus_{n\geq 0}A_{n}$ equipped with an action of $\mathfrak{S}_{n}$ on $A_{n}$ such that:
  1. (1)
    the multiplication map $A_{n}\otimes A_{m}\to A_{n+m}$ is $(\mathfrak{S}_{n}\times\mathfrak{S}_{m})$-equivariant (we use the standard embedding $\mathfrak{S}_{n}\times\mathfrak{S}_{m}\subset\mathfrak{S}_{n+m}$ for the action on $A_{n+m}$); and
  2. (2)
    given $x\in A_{n}$ and $y\in A_{m}$ we have $xy=\tau_{m,n}(yx)$, where $\tau_{m,n}\in\mathfrak{S}_{n+m}$ is defined by
    $\tau(i)=\begin{cases}i+n&\text{if }1\leq i\leq m,\\ i-m&\text{if }m+1\leq i\leq n+m.\end{cases}$
The last condition above is the “twisted commutativity” condition. ∎
Let $W$ be a $k$-vector space considered as a representation of $\mathfrak{S}_{1}$; we emphasize this by instead writing $W\langle 1\rangle$. The polynomial tca over $W\langle 1\rangle$ is the non-commutative polynomial ring $T(W)$ over $W$ with the $S_{n}$-action on $T(W)_{n}$ being the obvious one: on monomials it is given by $\sigma(w_{1}\cdots w_{n})=w_{{\sigma^{-1}(1)}}\cdots w_{{\sigma^{-1}(n)}}$.
2.3. Finite generation
Let $S$ be a $\mathbf{GL}$-algebra and $R$ be a tca. The $\mathbf{GL}$-algebra $S$ is finitely generated in degrees $\leq s$ if it is generated (as a $k$-algebra) by the $\mathbf{GL}$-orbit of finitely many homogeneous elements all of degree $\leq s$. Similarly, the tca $R$ is finitely generated in degrees $\leq s$ if there exist finitely many elements $r_{i}$ in various $R_{n_{i}}$ with $n_{i}\leq s$, such that the smallest subspace of $R$ containing the $r_{i}$ and stable under multiplication and the action of the symmetric groups, is $R$ itself. The $\mathbf{GL}$-algebra $\operatorname{Sym}(U)$ (resp. tca $\operatorname{Sym}(\{V_{n}\})$) is finitely generated if and only if $U$ is a finite-length polynomial representation (resp. $\sum_{n}\dim(V_{n})$ is finite).
We also consider module objects for $S$ in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$; this forms an abelian category $\operatorname{Mod}_{S}$. The $\mathbf{GL}$-algebra $S$ is weakly noetherian if $S$-submodules of $S$ satisfy the ascending chain condition, and it is noetherian if $\operatorname{Mod}_{S}$ is locally noetherian. These definitions extend analogously to the tca $R$.
2.4. The Schur functor
We have introduced two parallel concepts that, despite their apparent differences, are closely connected through the Schur functor. The Schur functor $\mathcal{S}\colon\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})\to\operatorname{Rep}(\mathfrak{S}_{*})$ is defined on objects by the formula
$\mathcal{S}(V)\coloneqq\{\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},V)\}_{n\in\mathbf{N}}$
which can be identified with the $(1^{n})$ weight space of $V$ with $n$ varying (see [Rei16, § 2.4] and references therein). The Schur functor is an equivalence of symmetric monoidal categories if and only if $p=0$ (see [SS12, §5.4]).
Lemma 2.3.
Assume $p>0$.
  1. (1)
    The Schur functor $\mathcal{S}$ is exact;
  2. (2)
    For a partition $\lambda$, we have $\mathcal{S}(L_{\lambda}(\mathbf{V}))\neq 0$ if and only if $\lambda$ is $p$-restricted.
  3. (3)
    For each partition $\lambda$, we have $\mathcal{S}(\mathbf{S}_{\lambda}(\mathbf{V}))=S^{\lambda}$, where $\mathbf{S}_{\lambda}(\mathbf{V})$ is the Schur module with highest weight $\lambda$ and $S^{\lambda}$ is the Specht module corresponding to $\lambda$.
Proof.
  1. (1)
    The exactness of $\mathcal{S}$ follows from, for example, the fact that $\mathbf{V}^{\otimes n}$ is a projective object in $\operatorname{Rep}^{\mathrm{pol}}(\mathbf{GL})$ (see [Rei16, Remark 2.10])
  2. (2)
    We have $\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},V)=(1^{n})$ weight space of $V$, and $(L_{\lambda}(\mathbf{V}))_{(1^{n})}\neq 0$ if and only if $\lambda$ is $p$-restricted (see [Rei16, § A.3]).
  3. (3)
    See [Rei16, § A.3]. ∎
We sketch the proof of the next well-known result, as it plays a crucial role in our method.
Proposition 2.4.
The Schur functor $\mathcal{S}$ is a symmetric monoidal functor.
Proof.
We first define a natural map $\mathcal{S}(U)\boxtimes\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$. Without loss of generality, we may assume $U$ and $W$ are finite length and homogeneous of degree $d$ and $e$ respectively. Recall that $\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes n},V)$ is naturally isomorphic to the $(1^{n})$ weight space of $V$ with the isomorphism given by $v\in V_{(1^{n})}\mapsto f_{v}$ where $f_{v}$ is the $\mathbf{GL}$-equivariant map which maps
$e_{1}\otimes e_{2}\otimes\ldots\otimes e_{n}\in\mathbf{V}^{\otimes n}\text{ to}v\in V.$
We define an $\mathfrak{S}_{d}\times\mathfrak{S}_{e}$-equivariant map from $\mathcal{S}(U)\times\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$ where $f_{u}\otimes f_{w}$ with $f_{u}\in\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes d},U)$ and $f_{w}\in\operatorname{Hom}_{\mathbf{GL}}(\mathbf{V}^{\otimes e},W)$ is mapped to $f_{u\otimes w}$. This extends to an $\mathfrak{S}_{d+e}$-equivariant map $\mathcal{S}(U)\boxtimes\mathcal{S}(W)\to\mathcal{S}(U\otimes W)$ and is natural in both variables. Furthermore, the domain and codomain (for fixed $U$ and $W$) both have dimension $\binom{d+e}{d}\dim(U_{1^{d}})\dim(W_{1^{e}})$ and by comparing weight spaces, we see that the map is surjective, so it is an isomorphism. The natural isomorphism defined clearly commutes with the symmetric structures, so $\mathcal{S}$ is a symmetric monoidal functor. ∎
This proposition implies that the Schur functor takes a $\mathbf{GL}$-algebras and twisted commutative algebras. We will mainly use the following concrete application.
Example 2.5.
Let $W$ be a finite-dimensional vector space. By Proposition 2.4, we have $\mathcal{S}(\operatorname{Sym}(W\otimes\mathbf{V}))\cong\operatorname{Sym}(\mathcal{S}(W\otimes\mathbf{V}))=\operatorname{Sym}(W\langle 1\rangle)=T(W)$. ∎
3. Finite group invariants
Throughout this section, we let $W$ be a representation of a group $G$, and $S$ and $R$ be the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})$ and tca $T(W)$ respectively.
3.1. The tca structure on $T(W)^{G}$
We briefly explain why $T(W)^{G}$ has the structure of a twisted commutative algebra. The tensor algebra $T(W)$ can be viewed as a tca with $W$ an $\mathfrak{S}_{1}$-representation. This is the free tca over $W$ as explained after Definition 2.2 so the action of $G$ on $W$ extends to a tca automorphism of $T(W)$ and thereby, the tca structure descends to $T(W)^{G}$.
Remark 3.1 (Freeness of the invariant ring).
For finite groups $G$, the invariant subring $\operatorname{Sym}(W)^{G}$ is a polynomial ring only if the action of $G$ is generated by pseudo-reflections. Remarkably, Kharchenko [Kha78] proved that $T(W)^{G}$ is always free as an associative $k$-algebra. The tca $T(W)^{G}$ is almost never a free tca as can be seen using growth rates: $T(W)^{G}_{n}$ grows exponentially with $n$ because $\dim(T(W)^{G}_{n})\leq\dim(W)^{n}$ but a free tca with at least one generator of degree $\geq 2$ grows superexponentially. ∎
For the next result, we freely use the fact that the Schur functor $\mathcal{S}$ applied to an object $A$ can be identified with the subspace spanned by all vectors of weight $(1^{n})$ for all $n\geq 0$ (see Section 2.4). For us, a flat weight is a weight of the form $(1^{n})$ for some $n\geq 1$, and a flat weight vector is a weight vector of flat weight.
Lemma 3.2.
The Schur functor induces an identification
$\mathcal{S}(S^{G})\cong R^{G}.$
Proof.
We identify $R$ with the subspace of $S$ spanned by all flat weight vectors. Under this identification, the action of $G$ on $R$ is just the induced action on this subspace. Similarly, we identify $\mathcal{S}(S^{G})$ with the subspace of $S^{G}$ spanned by all flat weight vectors. An element is in $\mathcal{S}(S^{G})$ if and only if it (1) is a flat weight vector and (2) is also invariant under the action of $G$. By (1), the element is in $R$ and by (2), the element is in $R^{G}$. The reverse containment is similar. ∎
Corollary 3.3.
Assume $S^{G}$ is finitely generated in degrees $\leq r$. Then $R^{G}$ is also finitely generated in degrees $\leq r$.
Proof.
If $S^{G}$ is finitely generated in degree $\leq r$, there exists a finite length $\mathbf{GL}$-representation $U$ of degree $\leq r$ with a surjection
$\operatorname{Sym}(U)\to S^{G}.$
Applying the Schur functor, and using Lemma 2.3(1) and Lemma 3.2, we get a surjection
$\mathcal{S}(\operatorname{Sym}(U))\to R^{G}.$
By Proposition 2.4, we have $\mathcal{S}(\operatorname{Sym}(U))\cong\operatorname{Sym}(\mathcal{S}(U))$. Furthermore, the Schur functor $\mathcal{S}$ takes finite length objects of degree $\leq r$ to finite length objects of degree $\leq r$, we obtain the result. ∎
For a $\mathbf{GL}$-algebra or tca $A$, we let $A_{+}$ be the ideal of positive degree elements.
Lemma 3.4.
Assume for all $n\in\mathbf{N}$, there exists a $p$-restricted partition $\lambda$ with $|\lambda|>n$ such that $L_{\lambda}(\mathbf{V})$ is an irreducible constituent of $S^{G}_{+}/(S^{G}_{+})^{2}$. Then $R^{G}$ is not a finitely generated tca.
Proof.
Using Lemma 3.2 and the exactness of the Schur functor, we have that $R^{G}_{+}/(R^{G}_{+})^{2}\cong\mathcal{S}(S^{G}/(S^{G}_{+})^{2})$. By Lemma 2.3(ii), we get $\mathcal{S}(L_{\lambda}(\mathbf{V}))\neq 0$ for $p$-restricted $\lambda$. Therefore $R^{G}_{+}/(R^{G}_{+})^{2}$ is supported in infinitely many degrees, so $R^{G}$ is not finitely generated. ∎
When $p>0$, we do not know of an example where $R^{G}$ is finitely generated but $S^{G}$ is not. By the above result, all irreducible constituents of $S^{G}_{+}/(S^{G}_{+})^{2}$ in sufficiently large degree must not be $p$-restricted.
3.2. Finite groups
In this subsection, we assume $G$ is a finite group.
The key result that enables us to prove Theorem 1.1 is a mild refinement of Richman’s remarkable work [Ric96] about the failure of Weyl’s polarization theorem for finite groups in the modular setting. Richman shows that if $\operatorname{char}(k)\mid\#G$, then $S^{G}$ contains generators of sufficiently large degree – this result is insufficient for us as the Schur functor kills elements of large multidegree.
Proposition 3.5.
Assume $G\subset GL(W)$ is a finite group with #$\#G$ divisible by $p$. For all $n\in\mathbf{N}$, there exists a $p$-restricted partition $\lambda$ with $|\lambda|>n$ such that $L_{\lambda}(\mathbf{V})$ is an irreducible constituent of $S^{G}_{+}/(S^{G}_{+})^{2}$.
Proof.
We explain how to obtain the result from [Ric96]. Set $m=n\frac{p^{\#G}-1}{p-1}$. In Proposition 9 of loc. cit., Richman shows that there exists a homogeneous element $f$ of degree $>\frac{m(p-1)}{p^{\#G}-1}=n$ which is nonzero in $S^{G}_{+}/(S^{G}_{+})^{2}$. This element $f$ contains a monomial $y$ that divides the monomial $y^{*}$ defined at the end of Page 36 of loc. cit.. The element $y^{*}$ is defined as the product $\prod(H_{i}(x_{i}))^{p-1}$ where $x_{i}$ is a linear form in $W$ and $H_{i}$ is the $i$-th inclusion $W\to W^{\oplus N}\subset W\otimes\mathbf{V}$. From this definition, we see that $y^{*}$ is a weight vector for $\mathbf{GL}$ of weight $\mu$ with $\mu_{i}=p-1$ for all $i$ with $\mu_{i}$ nonzero. So all nonzero components of the weight of $y$ is $<p$, and in turn, the same holds for $f$. Now, in a polynomial $\mathbf{GL}$-representation $U$ that has no $p$-restricted irreducible representation as a constituent, at least one component of each weight occurring in $U$ will be $\geq p$. Therefore, the $\mathbf{GL}$-representation generated by $f$ in $S^{G}/(S^{G}_{+})^{2}$ contains an irreducible $L_{\lambda}(\mathbf{V})$ with $\lambda$ $p$-restricted, as required. ∎
We can now prove the theorem on finite groups.
Proof of Theorem 1.1.
The tca $T(W)$ is noetherian [SS17]. If #$\#G\neq 0$ in $k$, then $T(W)^{G}$ is a direct summand of $T(W)$ using the Reynolds operator. Therefore, $T(W)^{G}$ is also noetherian. To obtain the degree bound for $T(W)^{G}$, note that $\operatorname{Sym}(W\otimes\mathbf{V})^{G}$ is generated in degrees $\leq|G|$ by [Fle00], and therefore so is $T(W)^{G}$ by Corollary 3.3.
In the case when $p\mid\#G$, the invariant ring $T(W)^{G}$ is not finitely generated by Proposition 3.5 and Lemma 3.4. ∎
In the next example, we explain how our ideas can be used to obtain explicit generators of invariant subrings, recovering results of [BDDK22, BDDK23].
Example 3.6.
Let $k$ be a field of characteristic $p>0$. Consider the $\mathfrak{S}_{p}$ action on $T(k^{p})=k\langle x_{1},x_{2},\ldots,x_{p}\rangle$. We identify $S=\operatorname{Sym}(k^{n}\otimes\mathbf{V})$ with $k[x_{i,j}|i\in[p],j\in\mathbf{N}]$. The invariant subring $S^{G}$ is not finitely generated [Ryd07]: the polynomials $f_{n}\coloneqq x_{1,1}x_{1,2}\ldots x_{1,n}+x_{2,1}x_{2,2}\ldots x_{2,n}+\ldots+x_{p,1}x_{p,2}\ldots x_{p,n}$ are nonzero in $S^{G}_{+}/(S^{G}_{+})^{2}$. Furthermore, up to scalars, the polynomials $f_{n}$ are the only nonzero elements of multidegree $(1^{n})$ in $S^{G}/(S^{G}_{+})^{2}$. The Schur functor maps $f_{n}$ to the element $p_{n}\coloneqq x_{1}^{n}+x_{2}^{n}+\ldots+x_{p}^{n}\in T(k^{p})$. So $\{p_{n}\}_{n}$ minimally generates $R^{G}$. ∎
4. Some results for invariants of classical groups
In this section alone, we further assume that our field $k$ is algebraically closed of characteristic $p>0$. Our exposition mostly follows [DM21].
Let $G$ be a connected split reductive group scheme over $\mathbf{Z}$. Let $W$ be a finite $\mathbf{Z}$-module with an algebraic action of $G$, i.e., $W$ is a comodule over the Hopf algebra $\mathbf{Z}[G]$. The base change $W_{k}$ is a rational representation of $G_{k}=G\times_{\mathbf{Z}}k$. We prove some results about the structure of $T(W_{k})^{G_{k}}$ as tca for specific representations $W$.
4.1. Universal good filtrations
We briefly recall the notion of a good filtration for representations of $G$ over $\mathbf{Z}$, assuming familiarity with the same notion over algebraically closed fields. A good reference is [Jan03, Part II], especially Appendix B.
We fix a Borel subgroup $B\subset G$ containing a maximal split torus $T$ of $G$. For a dominant integral weight $\lambda$, there exists a $G$-representation $\nabla(\lambda)$ such that the dual Weyl module of $G_{k}$ corresponding to $\lambda$ is isomorphic to the dual Weyl module $\nabla(\lambda)_{k}$ for every algebraically closed field $k$. We set $\nabla(\lambda)$ to be the dual Weyl module of $G$ corresponding to $\lambda$ over $\mathbf{Z}$. A $G$-representation has a universal good filtration if it has a filtration such that the successive quotients are isomorphic to dual Weyl modules $\nabla(\lambda)$ with $\lambda$ allowed to vary. Every $G$-representation which admits a universal good filtration is free as a $\mathbf{Z}$-module since the dual Weyl modules over $\mathbf{Z}$ are themselves free $\mathbf{Z}$-modules.
4.2. Results
We now prove a result that easily implies Theorem 1.2 since every classical group is connected and split reductive over $\mathbf{Z}$.
Theorem 4.1.
Assume $W$ is a finite $\mathbf{Z}$-module with an algebraic action of $G$ admitting a universal good filtration. For $p\gg 0$, the invariant ring $T(W_{k})^{G_{k}}$ is a finitely generated tca over $k$.
Proof.
By [DM21, Theorem 1.10(2)], the $\mathbf{GL}$-algebra $\operatorname{Sym}(W_{k}\otimes\mathbf{V})^{G_{k}}$ is finitely generated for $p\gg 0$. So $T(W_{k})^{G_{k}}$ is a finitely generated tca over $k$ by Corollary 3.3. ∎
Remark 4.2 (Small characteristics).
Consider the rank $n^{2}$ free $\mathbf{Z}$-module $W$ consisting of $n\times n$ square matrices endowed with the conjugation action of $\mathbf{GL}_{n}$. The representation $W$ has a good filtration over $\mathbf{Z}$ [DKZ02, Page 63]. However, the tca $T(W_{k})^{GL_{n}(k)}$ is not finitely generated if char $k\leq n$. Indeed, in Corollary 3.2 of loc. cit., the authors show that a minimal generating set of $\operatorname{Sym}(W\otimes\mathbf{V})^{GL_{n}(k)}$ consists of certain trace elements of weight $1^{m}$ for all $m>0$. The result now follows from Lemma 3.4 and Lemma 2.3(2). ∎
We now establish a result about the $\mathfrak{S}_{*}$-representation structure of $T(W_{k})^{G_{k}}$.
Theorem 4.3.
Assume $W$ is a $G$-representation over $\mathbf{Z}$ with a universal good filtration. For $p>\dim(W)$, the invariant ring $T(W_{k})^{G_{k}}$ has a Specht filtration.
Proof.
By [DM21, Corollary 6.8], when $\operatorname{char}(k)>\dim(W)$, the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})^{G_{k}}$ has a filtration with associated graded pieces being the Schur modules $\mathbf{S}_{\lambda}(\mathbf{V})$. The result now follows by combining (1) and (3) of Lemma 2.3. ∎
5. Gelfand–Kirillov dimension
Recall that $k$ is an infinite field of characteristic $p\geq 0$. Let $\mathcal{C}$ be a symmetric monoidal category over $k$ of Krull–Gabriel dimension zero, and $A$ be a finitely generated algebra object in $\mathcal{C}$ with $V\subset A$ being a finite length subobject of $A$ that generates it. Let $f_{A;V}(n)$ be the length of the image $A_{V,\leq n}$ of the canonical map $\bigoplus_{i\leq n}\operatorname{Sym}^{i}(V)\to A$.
Definition 5.1.
The categorical Gelfand–Kirillov dimension of $A$ with respect to $V$, denoted $\dim(A;V)$, is the smallest real number $r$ such that there exists a real number $c$ with $f_{A;V}(n)\leq cn^{r}$ for all $n$. If no such real number exists, we set $\dim(A;V)=\infty$.
As with the ordinary Gelfand–Kirillov dimensions, we have:
Proposition 5.2.
Assume $V$ and $W$ are two finite length subobjects of $A$ that generate it. Then $\dim(A;V)=\dim(A;W)$.
Proof.
It suffices to show that $\dim(A;V)\leq\dim(A;W)$. There exists $l>0$ such that $W\subset A_{V,\leq l}$, so we have $f_{A;W}(n)\leq f_{A;V}(nl)$ which in turn gives the required inequality. ∎
We set $\operatorname{GKdim}(A)=\dim(A;V)$ for any finite length generating subobject $V\subset A$; this is well-defined by the above proposition. We can similarly define $\operatorname{GKdim}$ for module objects for $A$ in $\mathcal{C}$. The next result, along with the example that follows, allows us to compute $\operatorname{GKdim}$ for invariant subrings of finite groups when $p=0$. Interestingly, it is finite.
Lemma 5.3.
Assume $A\to B$ is a map of algebra objects in $\mathcal{C}$ such that $B$ is finitely generated as an $A$-module and the map splits as a map of $A$-modules. Then $\operatorname{GKdim}(A)=\operatorname{GKdim}(B)$.
Proof.
This is similar to the proof of Proposition 5.2 so we skip it. ∎
Recall that the ordinary Gelfand–Kirillov dimension of noncommutative tensor algebras is infinite. We now compute $\operatorname{GKdim}(T(W))$ as a tca.
Example 5.4.
Let $W$ be an $n$-dimensional vector space. By Cauchy’s formula, the $\mathbf{GL}$-algebra $\operatorname{Sym}(W\otimes\mathbf{V})$ has a filtration with associated graded pieces $\mathbf{S}_{\lambda}(W)\otimes\mathbf{S}_{\lambda}(\mathbf{V})$ with $\lambda$ varying over all partitions. Applying the Schur functor and using Lemma 2.3, we get that $T(W)$ has a filtration with associated graded pieces $\mathbf{S}_{\lambda}(W)\otimes S^{\lambda}$. Therefore, we get
$f_{A;W}(N)=\sum_{|\lambda|\leq N}\dim(\mathbf{S}_{\lambda}(W))\operatorname{len}(S^{\lambda}).$
Recall that $\mathbf{S}_{\lambda}(W)$ vanishes if $\lambda$ has $>\dim(W)$ nonzero parts. For the sake of simplicity, we now assume $\operatorname{char}(k)=p=0$ so that $\operatorname{len}(S^{\lambda})=1$ in the above sum. We now estimate $f_{A;W}$: the number of partitions of $N$ with at most $n$ parts is $\approx cN^{n-1}$ with $c$ a constant for $N\gg n$; and the dimension of the Schur module $S_{\lambda}(W)$ with $|\lambda|\gg n$ and $\lambda_{n+1}=0$ is $\approx d|\lambda|^{\frac{n(n-1)}{2}}$ where $d$ is another constant, as can be seen using Weyl’s dimension formula. Therefore, $f_{A;W}(N)\approx\sum_{i=0}^{N}cdi^{n-1}i^{\frac{n(n-1)}{2}}$ which has dominant term $N^{\frac{n(n+1)}{2}}$, from which we get $\operatorname{GKdim}(T(W))=\binom{n+1}{2}$. ∎
It would be interesting to also compute $\operatorname{GKdim}$ of $T(W)$ when $p>0$. If $\dim(W)\in\{1,2\}$, then $\operatorname{GKdim}$ is independent of characteristic as $\operatorname{len}(S^{\lambda})$ for partitions with $\leq 2$ parts is at most $\approx\log(|\lambda|)$ (see Example 5.6).
We note that $\operatorname{GKdim}$ does not provide anything meaningful for polynomial tcas generated in degree $\geq 2$. We show it is infinite for the polynomial algebra of the trivial representation $\operatorname{triv}_{2}$ of $\mathfrak{S}_{2}$.
Example 5.5.
The $\mathbf{GL}$-algebra $\operatorname{Sym}(\operatorname{Sym}^{2}(\mathbf{V}))$ has a multiplicity-free filtration by Schur modules $\mathbf{S}_{2\lambda}(\mathbf{V})$ [Bof91], where $2\lambda$ is the partition $(2\lambda_{1},2\lambda_{2},\ldots)$. Applying the Schur functor, we see that the tca $A=\operatorname{Sym}(\operatorname{triv}_{2})$ has a filtration by the Specht modules $S_{2\lambda}$. Let $p(n)$ be the partition function. Then in characteristic zero, we have $f_{A;\operatorname{triv}_{2}}(n)=\sum_{i\leq n}p(i)$, and $f_{A;\operatorname{triv}_{2}}\geq\sum_{i\leq n}p(i)$ since $\operatorname{len}(S_{\lambda})\geq 1$. The partition function is subexponential so $\operatorname{GKdim}(A)=\infty$ in all characteristics. ∎
We finally compute $\operatorname{GKdim}$ for $T(k^{2})^{\mathbf{SL}_{2}(k)}$ with $k$ algebraically closed.
Example 5.6.
On the $\mathbf{GL}$-algebra side, the $\mathbf{SL}_{2}(k)$-invariant subring of $\operatorname{Sym}(\mathbf{V}\oplus\mathbf{V})$ is isomorphic to the quotient of $\operatorname{Sym}({\textstyle\bigwedge}^{2}(\mathbf{V}))$ by the ideal generated by the Plücker relations. By [Bof91, Theorem 2.7], this quotient ring has a multiplicity-free filtration by the Schur modules $\mathbf{S}_{(2^{n})}(\mathbf{V})$. Applying the Schur functor and using Lemmas 3.2 and 2.3(3), we see that $A\coloneqq T(k^{2})^{\mathbf{SL}_{2}(k)}$ has a multiplicity-free filtration by the Specht modules $S_{(2^{n})}$. If $\operatorname{char}(k)=p=0$, we immediately see that $\operatorname{GKdim}(A)=1$. If instead $p>0$, then we have $S^{\lambda}\cong S^{\lambda^{t}}\otimes\operatorname{sgn}$ where $\operatorname{sgn}$ is the sign representation of $\mathfrak{S}_{|\lambda|}$, so $\operatorname{len}(S_{(2^{n})})=\operatorname{len}(S_{(n,n)})$. By the work of James (see [Fay03, Theorem 2.5]), the length of $S_{(n,n)}$ is $\approx\lfloor\log_{p}(\frac{n+1}{2})\rfloor$. Combining the previous two facts, we get that $f_{A;A_{2}}(n+1)-f_{A;A_{2}}(n)\approx\lfloor\log_{p}(\frac{n+1}{2})\rfloor$ from which we also get $\operatorname{GKdim}(A)=1$. ∎
Acknowledgements.
The author thanks Mátyás Domokos, Martin Kassabov and Andrew Snowden for helpful discussions over email, and is especially grateful to Lucas Buzaglo for patiently explaining foundational results about noncommutative rings.